A student walks and jogs to college each day. she averages 5km/h walking and 9km/h jogging. The distancefrom home to college is 8km and she makes the trip in 1 hour. How far does the student jog?
I can't seem to get it set up right to get the correct calculation. I used 5km(X) +9 km(y) = 8km.





Answer by mananth(16949)

let her walk x km.
she will jog 8-x km
..
walking speed = 5km/hr
jogging speed = 9km/hr
..
distance to college 8km.
walking time + jogging time = total time
x/5+ 8-x /9 =1
9x+40-5x /45=1
4x+40=45
4x=5
x= 5/4 km
she walks 5/4 km
1.25 km
so she jogs 8-5/4
32-5 /4
27/4
6.75 km